The mean and standard deviation of serum iron values which are normally distributed for a random sample of 50 healthy men are 120 and 15 micrograms respectively. Find 95% confidence limits within the population mean. (z=1.96)


1
Expert's answer
2021-03-26T03:55:44-0400

Solution:

Given, "n=50,\overline{X}=120,\sigma=15"

For 95%, z=1.96

Now, confidence interval for "\mu=(\overline{X}-z(\sigma/\sqrt n), \overline{X}+z(\sigma/\sqrt n))"

"=(120-1.96(15/\sqrt {50}), 120+1.96(15/\sqrt {50}))
\\\approx(120-4.157,120+4.157)
\\=(115.843,124.157)"